# JWST PRIMER: A lack of outshining in four normal galaxies from the ALMA-CRISTAL Survey

### v1

@@ Abstract
We present a spatially resolved analysis of four star-forming galaxies at using data from theĀ JWST PRIMER and ALMA-CRISTAL surveys to probe the stellar and inter-stellar medium properties on the sub- scale. In the Ā JWST NIRCam imaging we find that the galaxies are composed of multiple clumps (between and ) separated by , with comparable morphologies and sizes in the rest-frame UV and optical. Using BAGPIPES to perform pixel-by-pixel SED fitting to theĀ JWST data we show that the SFR () and stellar mass ( ) derived from the resolved analysis are in close () agreement with those obtained by fitting the integrated photometry. In contrast to studies of lower-mass sources, we thus find a reduced impact of outshining of the older (more massive) stellar populations in these normal galaxies. OurĀ JWST analysis recovers bluer rest-frame UV slopes () and younger ages () than archival values. We find that the dust continuum from ALMA-CRISTAL seen in two of these galaxies correlates, as expected, with regions of redder rest-frame UV slopes and the SED-derived , as well as the peak in the stellar mass map. We compute the resolved āĀ relation, showing that the IRX is consistent with the local starburst attenuation curve and further demonstrating the presence of an inhomogeneous dust distribution within the galaxies. A comparison of the CRISTAL sources to those from the FirstLight zoom-in simulation of galaxies with the same Ā and SFR reveals similar age and colour gradients, suggesting that major mergers may be important in the formation of clumpy galaxies at this epoch.
@@
galaxies: high-redshift ā galaxies: ISM ā galaxies: irregular1 Introduction
@@
A full understanding of the formation and evolution of early galaxies requires spatially resolved information. An analysis of the emitted rest-frame UV and optical light provides a measure of the stellar properties at the time of observation, as well as the star formation history (SFH) of the galaxy by probing stars of different ages. The continuum and line emission in this wavelength range can also be impacted by (potentially inhomogeneous) dust attenuation (e.g.Ā Abdurroāuf etĀ al., 2023; Hashimoto etĀ al., 2023; Killi etĀ al., 2024) and/or metallicity gradients (e.g. identified through spectroscopic studies;Ā Tripodi etĀ al., 2024; Venturi etĀ al., 2024). The wavelength dependent morphology further provides information on the formation mechanism of high-redshift galaxies by highlighting any signatures of merging activity, for example through the identification of multiple discrete clumpsĀ (Bowler etĀ al., 2017; Barisic etĀ al., 2017; Hashimoto etĀ al., 2019), or established stellar disks through the detection of a centralised redder component attributed to older starsĀ (Ji etĀ al., 2023; Setton etĀ al., 2024; Fujimoto etĀ al., 2024).
@@
TheĀ Hubble Space Telescope (HST) has provided a wealth of information about the morphology and colour of galaxies up to very high redshifts (, e.g.Ā Oesch etĀ al., 2010; Holwerda etĀ al., 2015). However, at theĀ HST/Wide-Field Camera 3 (WFC3) imaging is limited to probing the rest-frame UV regime, where dust obscuration can play a significant role. It is therefore not clear if the clumpy morphology observed in many galaxies is a result of mergers or if these components are embedded in an underlying disk (e.g.Ā Barisic etĀ al., 2017). With the advent of imaging from theĀ James Webb Space Telescope (JWST) NIRCam instrument, it is possible for the first time to obtain sub-kpc scale imaging in the rest-frame optical for galaxies up to . These measurements provide a step change in the study of the rest-frame optical continuum and emission line properties at these redshifts, which were previously only inferred from relatively poor resolutionĀ Spitzer/IRAC photometry (e.g.Ā Stefanon etĀ al., 2023).
@@
The majority of known galaxies have been identified via the detection of the Lyman- break in the rest-frame UV, sometimes in combination with a narrow-band filter to pick up the Lyman- emission line itself (seeĀ Stark, 2016 for a review). The physical properties of the galaxies are then obtained through spectral energy distribution (SED) fitting of the observed optical and near-infrared (NIR) photometry. Due to the compact sizes of the majority of galaxies at , even inĀ JWST imaging the galaxies are marginally resolved (with half-light radii of ;Ā Morishita etĀ al., 2024; Ono etĀ al., 2024; Varadaraj etĀ al., 2024). Global properties such as the stellar mass and star-formation rate (SFR) are therefore derived from the integrated light of the source. While the biases that can occur in properties derived from SED fitting have been extensively discussed for galaxies at (e.gĀ Sorba & Sawicki, 2018; Leja etĀ al., 2019), it is not clear how discrepant resolved and integrated properties of galaxies might be at high redshift. For example, the reduced cosmic time available and the potential lack of dust attenuationĀ (Fudamoto etĀ al., 2020a; Bowler etĀ al., 2024) could in principle simplify the SED fitting process making globally derived properties robust. With the unique combination of sensitivity, spatial resolution and spectroscopic capabilities ofĀ JWST at , it is now possible to perform spatially resolved analyses that probe the rest-frame optical for the first time. This allows the testing of the previously obtained physical properties for a wide range of galaxies.
@@
UsingĀ JWST NIRCam imaging of five lensed ā galaxies,Ā GimĆ©nez-Arteaga etĀ al. (2023) found that when spatially resolving the galaxies, the rest-frame UV light was dominated by compact regions that show young ages, as revealed by strong [OIII] Ā emission lines. The resolved SED fitting revealed regions of the galaxies with substantial stellar mass, but less current star formation and older ages. While the light from these regions is measured in the integrated photometry, these areas of high stellar mass are āoutshoneā by the younger pockets of star formation, which then leads to an underestimate of the stellar mass when fitting to the integrated photometry by up to . A similar effect is seen in the strongly lensed galaxy nicknamed the āCosmic Grapesā introduced inĀ Fujimoto etĀ al. (2024). In this galaxy, Ā GimĆ©nez-Arteaga etĀ al. (2024) demonstrate the effect of outshining and the inability of a variety of typically assumed SFHs in reproducing the galaxy properties from the global photometry. These studies, based onĀ JWST-selected strongly lensed galaxies, demonstrate that outshining appears to be significant at in ātypicalā sources (lying on the galaxy main sequence). It might be expected that such effects would be reduced at higher stellar masses, where in general the galaxy is experiencing less extreme or dominant pockets of very young star formation. Indeed,Ā PĆ©rez-GonzĆ”lez etĀ al. (2023) find no systematic offset between stellar masses derived from a resolved and integrated SED fitting analysis using a sample of red galaxies at with . Further results at indicate a reduced effect of outshining at , recovering similar spatially resolved properties to those obtained from fitting the integrated/global photometryĀ (Nelson etĀ al., 2023; Setton etĀ al., 2024). The goal of this work is to connect these studies, which probe widely different mass ranges and are based on various sample selections, by determining the significance of outshining in ātypicalā main-sequence galaxies at .
@@
The ALMA-[CII] resolved ISM in star-forming galaxies with ALMA (CRISTAL: Herrera-Camus in prep.) large program aims to revolutionise studies of the rest-frame UV to FIR SEDs of galaxies, with high-resolution ( arcsec, equivalent to kpc at ) [CII]Ā and dust continuum observations in a sample of over 20 main-sequence galaxies. Previous observations of the FIR emission in high-redshift galaxies has shown that the stellar mass range of ā is key for understanding the build-up of dust at early times. As shown in samples of galaxies observed through various ALMA large programs, within this mass range the obscured SFR fraction of ā galaxies is Ā (Fudamoto etĀ al., 2020b; Algera etĀ al., 2023; Mitsuhashi etĀ al., 2024; Bowler etĀ al., 2024), with some sources showing Ā (Fudamoto etĀ al., 2021).
@@
CRISTAL has already revealed significant dust obscured star formationĀ (Mitsuhashi etĀ al., 2023) in typical ā galaxies. It has also discovered the complex highly-resolved [CII]Ā emission around two massive galaxies suggestive of a merging event (Posses etĀ al. 2024; Solimano etĀ al. 2024), along with the first significant sample of resolved rest-frame UV to FIR SED fitting in Li et al. (2024), dust temperature measurements in HZ10Ā (Villanueva etĀ al., 2024) and the measurement of [CII]Ā sizesĀ (Ikeda etĀ al., 2024).
@@
In this work we perform a spatially resolved SED fitting analysis of four normal ā galaxies with to investigate the effect of outshining in main-sequence star-forming galaxies and to understand the trends and potential biases in the previously obtained SED fitting properties of galaxies of this stellar mass. Our sample includes all of the sources that are part of the CRISTAL ALMA large program that overlap with the Public Release Imaging for Extragalactic Research (PRIMER;Ā Dunlop etĀ al., 2021)Ā JWST imaging, which observed the Cosmic Evolution Survey (COSMOS) field with eight NIRCam filters. These sources are extended on the scale of an arcsecond and are detected with a high signal-to-noise ratio (SNR), allowing a detailed spatially resolved analysis. The resolved parameters are then compared with high-spatial resolution ( arcsec) ALMA-CRISTAL dust continuum emission.
@@
The paper is structured as follows. We present the datasets we use in SectionĀ 2, followed by a description of our point spread function (PSF) homogenisation and resolved SED fitting analysis in SectionĀ 3. The spatially resolved SED fitting maps are presented in SectionĀ 4, along with a comparison of the resolved and integrated properties. In SectionĀ 5 we discuss our results and compute a resolved infrared-excess ( /)- relation for the sample (where is the rest-frame UV slope; ). We end with our conclusions in SectionĀ 7. We assume AB magnitudes throughout, and use the cosmological parameters and . At , 1 arcsec corresponds to a physical distance of .
@@ 2 Data and Sample
This study is based onĀ JWST imaging data from the PRIMER survey, which allows a pixel-by-pixel spatially resolved analysis of the four galaxies we consider. We then compare the resulting physical parameter maps to ALMA observations in Band 7 obtained as part of the CRISTAL large program. In this section we describe the main features of these datasets and the selection function of the four galaxies we study.
@@ 2.1 JWST PRIMER
PRIMER was a Cycle 1 program (PI: Dunlop, PID 1837) providing deep NIRCam imaging within the COSMOS and Ultra-Deep Survey (UDS) fields. We utilise the PRIMER internal V0.5 data release, which includes the first half of the dataset in the PRIMER-COSMOS field as observed in December 2022 and January 2023 (known as the COSMOS-2 observations). Further data, observed in April and May 2023 covered the second half, or COSMOS-1 observations, however the four sources in our analysis sit within the COSMOS-2 region and hence this additional data was not required for this analysis. The data was reduced using a custom version of theĀ JWST calibration pipeline, PENCIL (PRIMER enhanced NIRCam Image Processing Library; Magee in prep.), starting from the uncal.fits files. The reduction process removed excess striping, masked the residual snowball artefacts and removed wisp artefacts in the F150W and F200W filters. The astrometry was aligned to GAIA data release 3Ā (Gaia Collaboration etĀ al., 2023). The COSMOS-2 data we use consists of a mosaic of 19 NIRCam pointings across 8-band photometry (F090W, F115W, F150W, F200W, F277W, F356W, F410M, F444W) where the limiting depths are approximately (0.3 arcsec diameter aperture;Ā Donnan etĀ al., 2024). The mosaic has a pixel scale of 0.03 arcsec per pixel.
We verified the internal astrometric accuracy of the data using cut-outs of the size of pixels (or arcmin) centred on our targets. Using the sample of sources selected within these cut-outs, we determined that the relative image alignment has a precision of arcsec or less than one pixel. For our analysis we homogenised the PSF of the images using the WebbPSF models (Perrin etĀ al., 2012, 2014). Each filter image was matched to that of the F444W PSF using convolution kernels derived from the PSFs using the Wiener-Hunt deconvolution algorithm.
@@ 2.2 ALMA-CRISTAL large program
The four galaxies that we study in this work were observed as part of the ALMA-CRISTAL large program (2021.1.00280.L; PI: Rodrigo Herrera-Camus, see Herrera-Camus et al. in prep) in Cycle 8. The goal of CRISTAL was to provide high sensitivity and spatial resolution observations of the [CII] line and dust continuum emission in a subset of galaxies from the ALMA Large Program to Investigate C+ at Early Times (ALPINE; 2017.1.00428.L; PI: O. Le Fèvre; Le Fèvre et al., 2020; Faisst et al., 2020; Béthermin et al., 2020). Further constraints were placed on the sample selection so that the 19 CRISTAL galaxies had stellar masses derived by SED fitting using Le Phare (Arnouts et al., 1999; Ilbert et al., 2006) of , had HST data available and were within a factor of three of the evolving star-formation main sequence (see Herrera-Camus et al. in prep for details). The ALMA Band 7 observations are centred on the known frequency of the [CII] line as detected in ALPINE. Extended and compact configurations were used to provide a beam size of in the natural weighting, but with a maximum recoverable scale of . Two of the galaxies we study in this work, CRISTAL-11 and CRISTAL-13, are detected significantly () in the dust continuum first by ALPINE (Béthermin et al., 2020), and now by CRISTAL. CRISTAL-15 and CRISTAL-17 had negligible detection in the dust continuum, however we utilise upper limits on the  in our IRX analysis. The extraction of the flux and inferred luminosity are described in Mitsuhashi et al. (2023), and we utilise these values in our determination of the IRX, though we do not include the ALMA data in our SED fitting and therefore do not PSF match the NIRCam images to the ALMA resolution. For CRISTAL-11 the beam of the natural weighted data was arcsec, with a root-mean square (RMS) of . For CRISTAL-13 the beam of the natural weighted data was arcsec, with a RMS of . For CRISTAL-15 the beam of the natural weighted data was arcsec, with a RMS of . For CRISTAL-17 the beam of the natural weighted data was arcsec, with a RMS of .

### v2

@@ Abstract
We present a spatially resolved analysis of four star-forming galaxies at using data from the JWST PRIMER and ALMA-CRISTAL surveys to probe the stellar and inter-stellar medium properties on the sub- scale. In the JWST NIRCam imaging we find that the galaxies are composed of multiple clumps (between and ) separated by , with comparable morphologies and sizes in the rest-frame UV and optical. Using BAGPIPES to perform pixel-by-pixel SED fitting to the JWST data, we show that the SFR () and stellar mass ( ) derived from the resolved analysis are in close () agreement with those obtained by fitting the integrated photometry. In contrast to studies of lower-mass sources, we thus find a reduced impact of outshining of the older (more massive) stellar populations in these normal galaxies. Our JWST analysis recovers bluer rest-frame UV slopes () and younger ages () than archival values. We find that the dust continuum from ALMA-CRISTAL seen in two of these galaxies correlates, as expected, with regions of redder rest-frame UV slopes and the SED-derived , as well as the peak in the stellar mass map. We compute the resolved – relation, showing that the IRX is consistent with the local starburst attenuation curve and further demonstrating the presence of an inhomogeneous dust distribution within the galaxies. A comparison of the CRISTAL sources to those from the FirstLight zoom-in simulation of galaxies with the same and SFR reveals similar age and colour gradients, suggesting that major mergers may be important in the formation of clumpy galaxies at this epoch.
@@
galaxies: high-redshift – galaxies: ISM – galaxies: irregular1 Introduction
@@
A full understanding of the formation and evolution of early galaxies requires spatially resolved information. An analysis of the emitted rest-frame UV and optical light provides a measure of the stellar properties at the time of observation, as well as the star formation history (SFH) of the galaxy by probing stars of different ages. The continuum and line emission in this wavelength range can also be impacted by (potentially inhomogeneous) dust attenuation (e.g. Abdurro’uf et al., 2023; Hashimoto et al., 2023; Killi et al., 2024) and/or metallicity gradients (e.g. identified through spectroscopic studies; Tripodi et al., 2024; Venturi et al., 2024). The wavelength dependent morphology further provides information on the formation mechanism of high-redshift galaxies by highlighting any signatures of merging activity, for example through the identification of multiple discrete clumps (Bowler et al., 2017; Barisic et al., 2017; Hashimoto et al., 2019), or established stellar disks through the detection of a centralised redder component attributed to older stars (Ji et al., 2023; Setton et al., 2024; Fujimoto et al., 2024).
@@
The Hubble Space Telescope (HST) has provided a wealth of information about the morphology and colour of galaxies up to very high redshifts (, e.g. Oesch et al., 2010; Holwerda et al., 2015). However, at the HST/Wide-Field Camera 3 (WFC3) imaging is limited to probing the rest-frame UV regime, where dust obscuration can play a significant role. It is therefore not clear if the clumpy morphology observed in many galaxies is a result of mergers or if these components are embedded in an underlying disk (e.g. Barisic et al., 2017). With the advent of imaging from the James Webb Space Telescope (JWST) NIRCam instrument, it is possible for the first time to obtain sub-kpc scale imaging in the rest-frame optical for galaxies up to . These measurements provide a step change in the study of the rest-frame optical continuum and emission line properties at these redshifts, which were previously only inferred from relatively poor resolution Spitzer/IRAC photometry (e.g. Stefanon et al., 2023).
@@
The majority of known galaxies have been identified via the detection of the Lyman- break in the rest-frame UV, sometimes in combination with a narrow-band filter to pick up the Lyman- emission line itself (see Stark, 2016 for a review). The physical properties of the galaxies are then obtained through spectral energy distribution (SED) fitting of the observed optical and near-infrared (NIR) photometry. Due to the compact sizes of the majority of galaxies at , even in JWST imaging the galaxies are marginally resolved (with half-light radii of ; Morishita et al., 2024; Ono et al., 2024; Varadaraj et al., 2024). Global properties such as the stellar mass and star-formation rate (SFR) are therefore derived from the integrated light of the source. While the biases that can occur in properties derived from SED fitting have been extensively discussed for galaxies at (e.g Sorba & Sawicki, 2018; Leja et al., 2019), it is not clear how discrepant resolved and integrated properties of galaxies might be at high redshift. For example, the reduced cosmic time available and the potential lack of dust attenuation (Fudamoto et al., 2020a; Bowler et al., 2024) could in principle simplify the SED fitting process making globally derived properties robust. With the unique combination of sensitivity, spatial resolution and spectroscopic capabilities of JWST at , it is now possible to perform spatially resolved analyses that probe the rest-frame optical for the first time. This allows the testing of the previously obtained physical properties for a wide range of galaxies.
@@
Using JWST NIRCam imaging of five lensed – galaxies, Giménez-Arteaga et al. (2023) found that when spatially resolving the galaxies, the rest-frame UV light was dominated by compact regions that show young ages, as revealed by strong [OIII] emission lines. The resolved SED fitting revealed regions of the galaxies with substantial stellar mass, but less current star formation and older ages. While the light from these regions is measured in the integrated photometry, these areas of high stellar mass are ‘outshone’ by the younger pockets of star formation, which then leads to an underestimate of the stellar mass when fitting to the integrated photometry by up to . A similar effect is seen in the strongly lensed galaxy nicknamed the ‘Cosmic Grapes’ introduced in Fujimoto et al. (2024). In this galaxy, Giménez-Arteaga et al. (2024) demonstrate the effect of outshining and the inability of a variety of typically assumed SFHs in reproducing the galaxy properties from the global photometry. These studies, based on JWST-selected strongly lensed galaxies, demonstrate that outshining appears to be significant at in ‘typical’ sources (lying on the galaxy main sequence). It might be expected that such effects would be reduced at higher stellar masses, where in general the galaxy is experiencing less extreme or dominant pockets of very young star formation. Indeed, Pérez-González et al. (2023) find no systematic offset between stellar masses derived from a resolved and integrated SED fitting analysis using a sample of red galaxies at with . Further results at indicate a reduced effect of outshining at , recovering similar spatially resolved properties to those obtained from fitting the integrated/global photometry (Nelson et al., 2023; Setton et al., 2024). The goal of this work is to connect these studies, which probe widely different mass ranges and are based on various sample selections, by determining the significance of outshining in ‘typical’ main-sequence galaxies at .
@@
The ALMA-[CII] resolved ISM in star-forming galaxies with ALMA (CRISTAL: Herrera-Camus in prep.) large program aims to revolutionise studies of the rest-frame UV to FIR SEDs of galaxies, with high-resolution ( arcsec, equivalent to kpc at ) [CII] and dust continuum observations in a sample of over 20 main-sequence galaxies. Previous observations of the FIR emission in high-redshift galaxies has shown that the stellar mass range of – is key for understanding the build-up of dust at early times. As shown in samples of galaxies observed through various ALMA large programs, within this mass range the obscured SFR fraction of – galaxies is (Fudamoto et al., 2020b; Algera et al., 2023; Mitsuhashi et al., 2024; Bowler et al., 2024), with some sources showing (Fudamoto et al., 2021).
@@
CRISTAL has already revealed significant dust obscured star formation (Mitsuhashi et al., 2023) in typical – galaxies. It has also discovered the complex highly-resolved [CII] emission around two massive galaxies suggestive of a merging event (Posses et al. 2024; Solimano et al. 2024), along with the first significant sample of resolved rest-frame UV to FIR SED fitting in Li et al. (2024), dust temperature measurements in HZ10 (Villanueva et al., 2024) and the measurement of [CII] sizes (Ikeda et al., 2024).
@@
In this work we perform a spatially resolved SED fitting analysis of four normal – galaxies with to investigate the effect of outshining in main-sequence star-forming galaxies and to understand the trends and potential biases in the previously obtained SED fitting properties of galaxies of this stellar mass. Our sample includes all of the sources that are part of the CRISTAL ALMA large program that overlap with the Public Release Imaging for Extragalactic Research (PRIMER; Dunlop et al., 2021) JWST imaging, which observed the Cosmic Evolution Survey (COSMOS) field with eight NIRCam filters. These sources are extended on the scale of an arcsecond and are detected with a high signal-to-noise ratio (SNR), allowing a detailed spatially resolved analysis. The resolved parameters are then compared with high-spatial resolution ( arcsec) ALMA-CRISTAL dust continuum emission.
@@
The paper is structured as follows. We present the datasets we use in Section 2, followed by a description of our point spread function (PSF) homogenisation and resolved SED fitting analysis in Section 3. The spatially resolved SED fitting maps are presented in Section 4, along with a comparison of the resolved and integrated properties. In Section 5 we discuss our results and compute a resolved infrared-excess ( /)- relation for the sample (where is the rest-frame UV slope; ). We end with our conclusions in Section 7. We assume AB magnitudes throughout, and use the cosmological parameters and . At , 1 arcsec corresponds to a physical distance of .
@@ 2 Data and Sample
This study is based on JWST imaging data from the PRIMER survey, which allows a pixel-by-pixel spatially resolved analysis of the four galaxies we consider. We then compare the resulting physical parameter maps to ALMA observations in Band 7 obtained as part of the CRISTAL large program. In this section we describe the main features of these datasets and the selection function of the four galaxies we study.
@@ 2.1 JWST PRIMER
PRIMER was a Cycle 1 program (PI: Dunlop, PID 1837) providing deep NIRCam imaging within the COSMOS and Ultra-Deep Survey (UDS) fields. We utilise the PRIMER internal V0.5 data release, which includes the first half of the dataset in the PRIMER-COSMOS field as observed in December 2022 and January 2023 (known as the COSMOS-2 observations). Further data, observed in April and May 2023 covered the second half, or COSMOS-1 observations, however the four sources in our analysis sit within the COSMOS-2 region and hence this additional data was not required for this analysis. The data was reduced using a custom version of the JWST calibration pipeline, PENCIL (PRIMER enhanced NIRCam Image Processing Library; Magee in prep.), starting from the uncal.fits files. The reduction process removed excess striping, masked the residual snowball artefacts and removed wisp artefacts in the F150W and F200W filters. The astrometry was aligned to GAIA data release 3 (Gaia Collaboration et al., 2023). The COSMOS-2 data we use consists of a mosaic of 19 NIRCam pointings across 8-band photometry (F090W, F115W, F150W, F200W, F277W, F356W, F410M, F444W) where the limiting depths are approximately (0.3 arcsec diameter aperture; Donnan et al., 2024). The mosaic has a pixel scale of 0.03 arcsec per pixel.
We verified the internal astrometric accuracy of the data using cut-outs of the size of pixels (or arcmin) centred on our targets. Using the sample of sources selected within these cut-outs, we determined that the relative image alignment has a precision of arcsec or less than one pixel. For our analysis we homogenised the PSF of the images using the WebbPSF models (Perrin et al., 2012, 2014). Each filter image was matched to that of the F444W PSF using convolution kernels derived from the PSFs using the Wiener-Hunt deconvolution algorithm.
@@ 2.2 ALMA-CRISTAL large program
The four galaxies that we study in this work were observed as part of the ALMA-CRISTAL large program (2021.1.00280.L; PI: Rodrigo Herrera-Camus, see Herrera-Camus et al. in prep) in Cycle 8. The goal of CRISTAL was to provide high sensitivity and spatial resolution observations of the [CII] line and dust continuum emission in a subset of galaxies from the ALMA Large Program to Investigate C+ at Early Times (ALPINE; 2017.1.00428.L; PI: O. Le Fèvre; Le Fèvre et al., 2020; Faisst et al., 2020; Béthermin et al., 2020). Further constraints were placed on the sample selection so that the 19 CRISTAL galaxies had stellar masses derived by SED fitting using Le Phare (Arnouts et al., 1999; Ilbert et al., 2006) of , had HST data available and were within a factor of three of the evolving star-formation main sequence (see Herrera-Camus et al. in prep for details). The ALMA Band 7 observations are centred on the known frequency of the [CII] line as detected in ALPINE. Extended and compact configurations were used to provide a beam size of in the natural weighting, but with a maximum recoverable scale of . Two of the galaxies we study in this work, CRISTAL-11 and CRISTAL-13, are detected significantly () in the dust continuum first by ALPINE (Béthermin et al., 2020), and now by CRISTAL. CRISTAL-15 and CRISTAL-17 had negligible detection in the dust continuum, however we utilise upper limits on the in our IRX analysis. The extraction of the flux and inferred luminosity are described in Mitsuhashi et al. (2023), and we utilise these values in our determination of the IRX, though we do not include the ALMA data in our SED fitting and therefore do not PSF match the NIRCam images to the ALMA resolution. For CRISTAL-11 the beam of the natural weighted data was arcsec, with a root-mean square (RMS) of . For CRISTAL-13 the beam of the natural weighted data was arcsec, with a RMS of . For CRISTAL-15 the beam of the natural weighted data was arcsec, with a RMS of . For CRISTAL-17 the beam of the natural weighted data was arcsec, with a RMS of . The astrometric accuracy of the dust continuum images for CRISTAL-11 and CRISTAL-13, approximated using , are and , respectively.